A landscape contractor bids on jobs where he can make $3000 profit. The relative
frequency of getting 1, 2, 3, 4, 5,6,7,8 or 9 jobs per month are shown.
Number of jobs 1 2 3 4 5 6 7 8 9
Frequency 56 78 95 98 23 65 78 90 23
Find the contractor’s expected profit per month. Also find out the deviation of the landscape
contractor.

Answer by ikleyn(53765)


The total number of jobs  is  72 + 98 + 95 + 89 + 23 + 45 + 78 + 90 + 23 = 613.


The contractor’s expected profit per month =  = 


                                           = 300*4.514 = 13541 dollars (rounded).